OpenStudy (anonymous):

simplify (2sqrt5+3sqrt7)^2, show your work

OpenStudy (espex):

Foil and simplify.\[(2\sqrt{5}+3\sqrt{7})(2\sqrt{5}+3\sqrt{7})\]

OpenStudy (anonymous):

thank you @eSpeX